OpenStudy (rebecca1588):

What is the solution to the system of equations represented by the two equations? y=-2x+7 y=1/3x (__?_,__3_)

pooja195 (pooja195):

We know the value of y for both equations correct? so Equation 1 y=-2x+7 Equation 2 y=1/3x Using the substituion method we can solve for x and y we know what y equals so lets plug equation 1 into equation 2 -2x+7=1/3 can you solve for x?

OpenStudy (rebecca1588):

x=10/3?

pooja195 (pooja195):

good :)

pooja195 (pooja195):

Now plug the value of x into equation 1 y=-2(10/3)+7

OpenStudy (rebecca1588):

1/3 :)

pooja195 (pooja195):

Good :) so (10/3,1/3)

OpenStudy (rebecca1588):

but my answer only has 2 boxes?

pooja195 (pooja195):

type it in as a fraction

pooja195 (pooja195):

Wait these can be simplfied

pooja195 (pooja195):

y=1 and x=3

OpenStudy (rebecca1588):

Thanks Pooja :)

OpenStudy (rebecca1588):

So my answer is (1,3)? (just checking so I type it in right)

pooja195 (pooja195):

You flipped it (3,1)

pooja195 (pooja195):

np :)

OpenStudy (rebecca1588):

Thanks your the best! :D